OpenStudy (anonymous):

The table below shows the ticket rates for whale watching trips offered by Beluga Fun Tours. Age Ticket Price Under three years free Three to twelve years $34 Over twelve years $45 On a certain day, the company took 103 people on whale watching trips. There were 64 children aged twelve and under, of which some children were under three years. If x represents the number of children under three years, which equation can be used to find the value of x, where C represents the total amount of money collected from tickets that day? (64 – x)34 + (103 – 64) ∙ 45 = C (103 + x)34 + (103

OpenStudy (anonymous):

he table below shows the ticket rates for whale watching trips offered by Beluga Fun Tours. Age Ticket Price Under three years free Three to twelve years $34 Over twelve years $45 On a certain day, the company took 103 people on whale watching trips. There were 64 children aged twelve and under, of which some children were under three years. If x represents the number of children under three years, which equation can be used to find the value of x, where C represents the total amount of money collected from tickets that day? (64 – x)34 + (103 – 64) ∙ 45 = C (103 + x)34 + (103 – (64 + x)) ∙ 45 = C (103 - x)34 + (103 – 64) ∙ 45 = C (64 + x)34 + (103 – (64 + x)) ∙ 45 = C

OpenStudy (ganpat):

x = under three= y = 3 -12 = z= 12+ = x + y = 64 ----? y = 64 - x 103 = x + y + z ---> 103 = 64 + z ---> z = 39 C = 34y + 45z C = 34 ( 64 - x ) + 45 * 39 C = 34 ( 64 - x ) + 1755... or option A ..

OpenStudy (anonymous):

so is the answer a